Malitz Inc. recently hired you as a consultant to estimate the company’s WACC. You have obtained the following information.

– Malitz’s noncallable bonds mature in 25 years, have an 8.00% annual coupon, a par value of $1,000, and a market price of $1,075.00.

– The company’s tax rate is 40%.

– The risk-free rate is 4.50%, the market risk premium is 5.50%, and the stock’s beta is 1.20.

– The target capital structure consists of 35% debt and the balance as common equity. Malitz uses the CAPM to estimate the cost of equity, and it does not expect to issue any new common stock. What is its WACC?

a.

7.51%

b.

7.90%

c.

8.32%

d.

8.76%

A firms cost of capital is often a reflection of its activities and funding needs. Consider the case of Wizard Company, and answer the following questions: Wizard Co. currently has only a real estate division and uses only equity capital; however, it is considering creating consulting and distribution divisions. Its beta is currently 1.3. The risk-free rate is 3.1%, and the market-risk premium is 6.2% O 7.13% O 11.16% O 3.10% O 6.20% This means that the firms real estate division will have a cost of capital of: The consulting division is expected to have a beta of 2.2, because it will be riskier than the firms real estate division O O 18.09% 19.24% 17.69% 16.74% This means that the firms consulting division will have a cost of capital of: The distribution division will have less risk than the firms real estate division, so its beta is expected to be 0.5. O 18.79% O 17.49% O 6.20% O 18.69% This means that the distribution divisions cost of capital will be: Wizard Co. expects 70% of its total value to end up in the real estate division, 20% in the consulting division, and 10% in the distribution division O 11.78% O 16.53% O 13.08% O 14.63% Based on this information, what rate of return should its investors require once it opens the new divisions?
Answer #1
Amount weight cost weight*cost
equity 0.65 0.6500 11.10% 0.0722
debt 0.35 0.3500 4.40% 0.0154

WACC = 8.76%

real estate division cost of capital = 3.1 + 1.3*6.2 = 11.16%

consulting division cost of capitl = 3.1 + 2.2*6.2 = 16.74%

distribution division = 3.1 + 0.5*6.2 = 6.20%

rate of return to investors = 0.7*11.16 + 0.2*16.72 + 0.1*6.2 = 11.78%